OpenStudy (anonymous):

2. A snack of an apple, a brownie, and a cheese stick has 270 calories. Two brownies and an apple have 290 calories. Two apples and three cheese sticks have 420 calories. How many calories are in each item?

OpenStudy (anonymous):

Basically, I just need help setting the problem up, and someone to check my answer at the end. I know I have to set it up as a system.

OpenStudy (anonymous):

what answer did you get ?

OpenStudy (anonymous):

I haven't gotten one, I'm still trying to set up the problem..

OpenStudy (anonymous):

I have a+b+c=270 2b+a=290 2a+3c=420 Is this set up correctly?

OpenStudy (anonymous):

yeah you set it up correctly... any idea on how to carry on ?

OpenStudy (anonymous):

My last equation is not the same as yours. Is yours right?

OpenStudy (anonymous):

But yes I think I do know how to go from here, but I may need someone to check my final answer. :)

OpenStudy (anonymous):

Yeah yous is right i miss some when writing equation ... sorry :)

OpenStudy (anonymous):

we ll check both our answers .. :D

OpenStudy (anonymous):

A = 90 Calories

OpenStudy (anonymous):

got that ?

OpenStudy (anonymous):

B = 100 Calories and C = 80 Calories

OpenStudy (anonymous):

Can you show me how you got that? I'm having problems :(

OpenStudy (anonymous):

\[From \space Equation \space 3 : C = \frac{ 420 - 2A }{ 3 }\] \[From \space Equation \space 2 : B \space = \frac{ 290 - A }{ 2 }\] \[From \space Equation \space 1 : A \space + B \space + C \space = \space 270\] \[A + \frac{ 290 - A }{ 2 } + \frac{ 420 -2A }{ 3 } =270 \]

OpenStudy (anonymous):

can you complete it now ? :D

OpenStudy (anonymous):

Yes, thank you very much!! :)

OpenStudy (anonymous):

no problem :D Once you get the value of A Just substitute in B = ___________ and C = ________ to have their value